OpenStudy (anonymous):

What is the radius of the circle whose equation is x2 + y2 = 8? 2squard2 64

OpenStudy (anonymous):

or 2

OpenStudy (experimentx):

the first one

OpenStudy (hoblos):

sqrt8 = sqrt(4*2) = 2sqrt2

OpenStudy (anonymous):

The equation of the circle is \[(x-a)^2+(y-b)^2=r^2\] where \[(a,b)\] is the center of the circle.

OpenStudy (anonymous):

r is the radius.